Status Report

Hubble Space Telescope Daily Report #2742 11/03/00

By SpaceRef Editor
November 3, 2000
Filed under

PERIOD COVERED:  0000Z (UTC) 11/02/00 – 0000Z (UTC) 11/03/00
Daily Status Report as of 308/0000Z
    1.1 Completed Three Sets of STIS/CCD 8837 (CCD Dark Monitor-Part 1)
        The Space Telescope Imaging Spectrograph (CCD) was used to monitor the darks.  The proposal completed with no reported problems.
    1.2 Completed STIS/CCD 8631 (Bright Quasar Close Lensing Search II)
        The Space Telescope Imaging Spectrograph (CCD) was used to expand the Cycle 8 second generation HST snapshot survey of bright quasars, optimized to find lenses with component image separations < 1".  The observations completed nominally.
    1.3 Completed STIS/CCD/MA2 8111 (CS 22892–052: A Rosetta Star for the Age and Early History of the Galaxy)
        The Space Telescope Imaging Spectrograph (CCD and MA2) was used to make observations of the "Rosetta" star CS 22892.  These observations should provide significant new information about the early history of the galaxy, specifically the nature of the earliest generations of stars and the types of nucleosynthetic processes that occurred in those stars.  The observations were executed as scheduled, and no anomalies were noted.
    1.4 Completed STIS/MA2 8843 (Cycle 9 MAMA Dark Measurements)
        The Space Telescope Imaging Spectrograph (MA2) was used to perform the routine monitoring of the MAMA detector dark noise.  The proposal completed nominally.
    1.5 Completed STIS/CCD/MA1 8103 (A Global Picture of White Dwarfs in Cataclysmic Variables)
        The Space Telescope Imaging Spectrograph (CCD and MA1) was used to make ultraviolet spectrographic observations of the white dwarf star that is located in the cataclysmic variable star system known as EG-CNC.  The observations were executed as scheduled with no problems.
    1.6 Completed STIS/CCD 8838 (Bias Monitor – Part 1)
        The Space Telescope Imaging Spectrograph (CCD) was used to monitor the bias in the 1×1, 1×2, 2×1, and 2×2 bin settings at gain=1, and 1×1 at gain = 4, to build up high-S/N superbiases and track the evolution of hot columns.  The proposal completed nominally.
    1.7 Completed Two Sets of STIS/CCD 8562 (Probing the Large Scale Structure: Cosmic Shear Observations)
        The Space Telescope Imaging Spectrograph (CCD) was used to probe the distortion of light bundles from distant galaxies, looking at the statistical properties of the intervening inhomogeneous {dark} matter distribution.  The proposal completed nominally.
The following scheduled WF/PC-2 observations did not execute due to the WF/PC-2 safing:
WF/PC-2 8822 (Decontaminations and Associated Observations Pt. 1/4) WF/PC-2 8326 (Colliding Stellar Winds and Proto-Planetary Dynamics) Four Sets of WF/PC-2 8826 (Cycle 9 Supplemental Darks pt 1)
    2.1 Guide Star Acquisitions:
      Scheduled Acquisitions: 9
                Successful: 9
    Scheduled Re-acquisitions: 7
                Successful: 7
    2.2 FHST Updates:
                  Scheduled: 20
                Successful: 20
    2.3 Operations Notes:
        WF/PC-2 remains in safe.  Recovery is not planned until next week.
        Using ROP SR-1A, the SSR EDAC error counter was cleared three times.
        One seven occasions, SI C&DH errors were cleared in accordance with ROP NS-5.
        Per ROP NS-3, the NSSC-1 status buffer was dumped at 307/1817Z.
        A TTR was written for a required re-transmit at 308/0441Z during a NSSC-1 uplink.
        SSA transmitter #2 was turned on at 308/0715Z and turned off at 308/0736Z, using ROP IC-2.
        Continuation of normal science observations and calibrations.

SpaceRef staff editor.